During the COVID-19 pandemic, the hospital emergency department at a very busy hospital receives, on average, 23.6 patient per day. The standard deviation of inter-arrival time between two consecutive patients arriving is 6.9 minutes. What is the coefficient of variation for the inter-arrival times (times between two consecutive arrivals) for patients arriving (received at) at the hospital’s emergency department? Note that, the emergency service of this hospital is open 24/7, and one day is equivalent to 24 hours.
Answer #1

Average time of per patient arrival (u) = 23.6 minutes

Standard deviation of inter arrival time between two patients (sd) = 6.9 minutes

Hence, Coefficient of variation for the inter arrival times = (sd / u)*100

                                                         = (6.9/ 23.6)*100

                                                         = 0.2924*100

                                                         = 29.24% (rounded off to 2 decimal places)